5. Express as a double integral the area enclosed by one loop of the curve 𝑟 = 3 cos 2𝜃 and



evaluate the integral

1
Expert's answer
2022-03-25T06:11:50-0400

One loop of the curve is presented on the image below:



where "\theta\in[-\frac{\pi}{4},\frac{\pi}{4}]". "r\in[0,3]."

Remind that for the polar coordinates: "x=r\,cos\,\theta" and "y=r\,sin\,\theta". "x\in[0,3\frac{\sqrt{2}}{2}]", "y\in[-3\frac{\sqrt{2}}{2},3\frac{\sqrt{2}}{2}]" It implies "x^2+y^2=r^2". The trigonometric formula for "cos\,2\theta" has the form: "cos\,2\theta=2\,cos^2\theta-1=\frac{2x^2}{r^2}-1=\frac{2x^2}{x^2+y^2}-1". Thus, the formula of the curve takes the form: "\sqrt{x^2+y^2}=\frac{2x^2}{x^2+y^2}-1". Simplifying the latter we receive: "(x^2+y^2)^{3\/2}=x^2-y^2". The latter implies "(x^2+y^2)^3=(x^2-y^2)^2" . The double integral for the area takes the form: "S=\int\int_{D}dxdy", where "D" is the area enclosed by the curve on the picture. I.e., "(x^2+y^2)^3=(x^2-y^2)^2", where "x\in[0,3\frac{\sqrt{2}}{2}]", "y\in[-3\frac{\sqrt{2}}{2},3\frac{\sqrt{2}}{2}]". However, it is easier to compute such an area in polar coordinates. The formula is: "S=\frac{1}{2}\int_{-\pi/4}^{\pi/4}r^2d\theta=\frac{1}{2}\int_{-\pi/4}^{\pi/4}9cos^2(2\theta)d\theta=\frac{9}{2}\int_{-\pi/4}^{\pi/4}\frac{1+cos(4\theta)}{2}d\theta=\frac{9}{4}(\theta+\frac{sin(4\theta)}{4})|_{-\pi/4}^{\pi/4}=\frac{9}{8}\pi"

The answer is "\frac{9}{8}\pi."